A 11.8 kg bowling ball traveling at 13.86 m/s strikes a 3.0 kg bowling pin. The impact slows the bowling ball to 7.60 m/s.



Determine the impulse given to the pin.
1
Expert's answer
2020-01-31T09:56:20-0500
"P=M(v_f-v_i)"

"P=11.8(13.86-7.6)=73.9\frac{kg\cdot m}{s}"